Complete the following table where indicated?

Source

SS

df

MS

F

Between

117.17

(b)_____

(c)_____

(e)_____

Within

(a)_____

9

(d)_____

Total

177.67

11


      Analyze the data from an independent groups design using ANOVA.

      a.      What is the value of dfwithin?
      b.      What is the value of dfbetween?
      c.      What is the value of Fobt?
      d.      Would you reject H0 using a = 0.05?
Answer #1

a) SSwithin = SStotal - SSbetween = 177.67 - 117.17

=> SSwithin = 60.5

b) dfbetween = dftotal - dfwithin = 11 - 9 = 2

=> dfbetween = 2

c)

MSbetween= SSbetween / dfbetween= 117.17/2 =58.585

MSwithin = SSwithin / dfwithin = 60.5/9 = 6.7222

Fobt = MSbetween / MSwithin = 58.582 / 6.7222 = 8.7152

=> Fobt = 8.72

d) F critical value at a = 0.05 with df1 = 2 and df2 = 9 is,

Fcrit = 4.256

Since, Fobt = 8.72 > Fcrit = 4.256

=> We reject H0